OpenStudy (anonymous):

Please help. I've been here for hours and no one has helped me I will medal, fan and testimonial

OpenStudy (michele_laino):

where is your question, please?

OpenStudy (anonymous):

its actually a download and openstudy wont let me upload it, so do you have an email that i can send it to?

OpenStudy (michele_laino):

can you make a screenshot of your question?

OpenStudy (anonymous):

oh nevermind it works :)

OpenStudy (anonymous):

its a portfolio and I really need help and its due today

OpenStudy (michele_laino):

question#1 the domain is the set of all values of the variable x

OpenStudy (anonymous):

got it, the range?

OpenStudy (anonymous):

doesnt it represents the people that showed up?

OpenStudy (michele_laino):

the range is the set of all values of the y-variable

OpenStudy (anonymous):

got it, can you help me out with #3?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Thanks so much !

OpenStudy (anonymous):

I dont know how to graph this

OpenStudy (michele_laino):

in order to get your graph, we have to use the table above, namely we have to draw the subsequent points: (1,8), (1,35), (3,62), (4,89)

OpenStudy (michele_laino):

oops..(2,35)

OpenStudy (anonymous):

Okay, can you draw it out I dont really understand

OpenStudy (anonymous):

still there?

OpenStudy (michele_laino):

here is the requested graph: 

OpenStudy (michele_laino):

I think that your graph is a line

OpenStudy (michele_laino):

the main feature, is that your line passes at the origin of the coordinate system

OpenStudy (anonymous):

okay so i would right "1 2 3 4" under my graph?

OpenStudy (michele_laino):

we can write its formula, making this conjecture: y=A*x+B

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

okay (: I'm ready for #4

OpenStudy (michele_laino):

your graph, as I wrote before, passes at the origin of the coordinate system

OpenStudy (michele_laino):

I think that it is the one of the features. Another feature, is that it is a straight line

OpenStudy (anonymous):

OH, but what about #3

OpenStudy (michele_laino):

it is a straight line, since we can see it from the same graph, even if it is an experimental graph

OpenStudy (anonymous):

thank you! i already have # 5 so i just need # 6-12

OpenStudy (michele_laino):

we have to start from this formula: y=A*x+B, where A, and B are two coefficient to be determined

OpenStudy (michele_laino):

coefficients*

OpenStudy (michele_laino):

if we set x=1 and y=8, I get: 8=A+B